OpenStudy (anonymous):

PLEASE HELP!!! Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. y = x2 + 4x − __

OpenStudy (anonymous):

substitute the values -16= (-2)^2 + 4(-2) - c -16 = 4 - 8 - c c = 4 - 8 +16 c = 12 y = x2 + 4x - 12

OpenStudy (anonymous):

thank you so much! i thought the answer was -20